A small business plans to target their sales to the upper 60% of income families. A study states that the average income per family is 62000. And the standard devation is 5000 find the cutoff target income per family. Please draw everything out so I can understand
Answer #1

Ans:

Given that

mean=62000

standard deviation=5000

P(Z>z)=0.60

P(Z<=z)=1-0.60=0.40

z=normsinv(0.40)

z=-0.25

Cut off value=62000-0.25*5000=60750